OpenStudy (anonymous):

would the derivative of (x^2/a^2 + y^2/b^2=1) be x^2+y^2=0

OpenStudy (anonymous):

i am doing implicit differentiation

OpenStudy (anonymous):

derivative with respect to \(x\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

start with \[\frac{2x}{a^2}+\frac{2y}{b^2}y'=0\] then solve for \(y'\)

OpenStudy (anonymous):

ok but my professor said we can use the constant multiple rule. how would that apply

OpenStudy (anonymous):

i am not sure what "constant multiple" rule means, but if it means \((cf)'=cf'\) then that is why the \(a^2\) and \(b^2\) stay in the denominator since \[\frac{x^2}{a^2}=\frac{1}{a^2}\times x^2\]

OpenStudy (anonymous):

rule is [cf(x)]'=cf'

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

what do i do after i have solved for y' but still have the a^2 and b^2

OpenStudy (anonymous):

yes they are still there they are just number, like 9 and 25 for example

OpenStudy (anonymous):

subtract \(\frac{2x}{a^2}\) from both sides, then multiply both sides by \(\frac{b^2}{2y}\)

OpenStudy (anonymous):

yes i know i got the answer. but what am i supposed to do with the constants?